OpenStudy (anonymous):

Choose two axioms that allow 6 + (x + 5) to be written x + 11. commutative - addition distributive associative - multiplication symmetric commutative - multiplication associative - addition identity - addition

OpenStudy (anonymous):

@Mimi_x3

OpenStudy (mimi_x3):

Well for starters we are distributing the plus sign over to x and 5

OpenStudy (haleyelizabeth2017):

So that pretty much gives you the answer ;)

OpenStudy (mimi_x3):

We need 2 axioms

OpenStudy (haleyelizabeth2017):

true....

OpenStudy (mimi_x3):

Ok addition associative laws are as follows (a+b)+c=a+(b+c) DO you see how this come into play?

OpenStudy (mimi_x3):

So 6+(x+5) can be rearranged to x+(6+5)=x+11